40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 286 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  2713   Thu Mar 25 09:07:08 2010 steveUpdatePEMGuralp2 x problem is back

Quote:

I went and double-checked and aligned the styrofoam cooler at ~5:00 UTC. It was fine, but we really need a better huddling box. Where's that granite anyway?

Here's the new Huddle Test output. This time I show the X-axis since there's some coherence now below 0.1 Hz.

You'll also notice that the Wiener filter is now beating the FD subtraction. This happened when I increased the # of taps to 8000. Looks like the noise keeps getting lower as I increase the number of taps, but this is really a kind of cheat if you think about it carefully.

 The same thing happening again.  The intermittent offset upstream of the seismometer that never got fixed.

The granite plate and ball bearings are in. I will place seismometers on it.

  2712   Wed Mar 24 15:59:59 2010 kiwamu, mattUpdateGreen Lockingleave PLL locked

Matt checked it in this morning and he found it's been locked during the night.

 

 

  2711   Wed Mar 24 14:57:21 2010 AlbertoUpdate40m UpgradingREFL11 upgraded

 

 Hartmut suggested a possible explanation for the way the electronics transfer function starts picking up at ~50MHz. He said that the 10KOhm resistance in series with the Test Input connector of the box might have some parasitic capacitance that at high frequency lowers the input impedance.

Although Hartmut also admitted that considering the high frequency at which the effect is observed, anything can be happening with the electronics inside of the box.

  2710   Wed Mar 24 14:52:02 2010 AlbertoUpdateGreen Lockingtwo NPRO PLL

Quote:

Now some pedestals, mirrors and lenses are left on the PSL table, since we are on the middle way to construct a PLL setup which employs two NPROs instead of use of PSL laser.

So Please Don't steal any of them.

 Can I please get the network analyzer back?

  2709   Wed Mar 24 12:40:25 2010 daisukeConfigurationGeneralPeriscope for green laser delivery from the BSC to PSL table

The periscope design for beam elevation of the green beams is posted. The design for the 90 deg steering version is also coming.

(2010-03-29: update drawings by daisuke)

90deg version: http://nodus.ligo.caltech.edu:8080/40m/2725

40m_periscope.png

  2708   Wed Mar 24 12:38:17 2010 HartmutConfigurationGreen LockingBroadband PD for green PLL

Modified one of the PD assemblies carrying a large SI-Diode (~10mm diameter).

Removed elements used for resonant operation and changed PD readout to transimpedance

configuration. The opamp is a CLC409 with 240 Ohm feedback (i.e. transimpedance) resistor.

To prevent noise peaking at very high frequencies and get some decoupling of the PD,

I added a small series resistor in line with the PD and the inverting opamp input.

It was chosen as 13 Ohm, and still allows for operation up to ~100MHz.

Perhaps it could be smaller, but much more bandwith seems not possible with this opamp anyway.

Changes are marked in the schematic, and I list affected components here.

(Numbers refer to version 'PD327.SCH' from 30-April-1997):

-removed L4

-connected L3 (now open pad) via 100 Ohm to RF opamp output. This restores the DC sognal output.

-removed c17

-connected pin 3 of opamp via 25 Ohm to GND

-connected kathode of PD via 13 Ohm to pin 2 of opamp

-removed L6, C26, L5, C18, and C27

-shorted C27 pad to get signal to the RF output

 

Measured the optical TF with the test laser setup.

(Note that this is at 1064nm, although the PD is meant to work with green light at 532nm!)

Essentially it looks usable out to 100MHz, where the gain dropped only by about

6dB compared to 10MHz.

Beyond 100MHz the TF falls pretty steeply then, probably dominated by the opamp.

 

The maximal bias used is -150V.

If the bias is 'reduced' from -150V to -50V, the response goes down by 4dB at 10MHz and

by 9dB at 100MHz.

 The average output was 30mV at the RF output, corresponding to 60mV at the opamp output (50Ohm divider chain).

With 240 Ohm transimpedance this yields 250µA photo-current used for these transfer functions.

SiAmpl.png

 

SiPhase.png

 

 

  2707   Wed Mar 24 04:22:51 2010 kiwamu, mattUpdateGreen Lockingtwo NPRO PLL

Now some pedestals, mirrors and lenses are left on the PSL table, since we are on the middle way to construct a PLL setup which employs two NPROs instead of use of PSL laser.

So Please Don't steal any of them.

  2706   Wed Mar 24 03:58:18 2010 kiwamu, matt, kojiUpdateGreen Lockingleave PLL locked

We are leaving the PLL as it is locked in order to see the long term stability. And we will check the results in early morning of tomorrow.

DO NOT disturb our PLL !!

  


(what we did)

After Mott left, Matt and I started to put feedback signals to the temperature control of NPRO.

During doing some trials Matt found that NPRO temperature control input has an input resistance of 10kOhm.

Then we put a flat filter ( just a voltage divider made by a resistor of ~300kOhm and the input impedance ) with a gain of 0.03 for the temperature control to inject a relatively small signal, and we could get the lock with the pzt feedback and it.

In addition, to obtain more stable lock we then also tried to put an integration filter which can have more gain below 0.5Hz.

After some iterations we finally made a right filter which is shown in the attached picture and succeeded in obtaining stable lock.

 

 

 

  2705   Wed Mar 24 02:06:24 2010 KojiUpdateIOOvac envelope has to be sealed as antproof for overnight

Matt and Koji:

We closed the light doors of the chambers.

Quote:

Roger.

Quote:

 This seal is good for daily use- operation only. The IFO has to be sealed  with light metal doors every night so ants and other bugs can not find their way in.

 

 

  2704   Tue Mar 23 22:46:43 2010 AlbertoUpdate40m UpgradingREFL11 upgraded
I modified REFL11 according to the changes lsited in this schematic (see wiki  / Upgrade 09 / RF System / Upgraded RF Photodiodes ).
I tuned it to be resonant at 11.06MHz and to have a notch at 22.12MHz.
These are the transfer functions that I measured compared with what I expected from the LISO model.

2010-03-23_REFL11_model_to_meas_comparison.png

The electronics transfer function is measured directily between the "Test Input" and the "RF Out" connector of the box. the optical transfer function is measured by means of a AM laser (the "Jenne laser") modulated by the network analyzer.
The AM laser's current was set at 20.0mA and the DC output of the photodiode box read about 40mV.
The LISO model has a different overall gain compared to the measured one, probably because it does not include the rest of the parts of the circuit other than the RF out path.

I spent some time trying to understand how touching the metal cage inside or bending the PCB board affected the photodiode response. It turned out that there was some weak soldering of one of the inductors.

  2703   Tue Mar 23 18:44:46 2010 MottUpdateABSLPLL reconstructed

 

 After realigning and getting the lock today, I tried to add in the SR785 to measure the transfer function.  As soon as I turn on the piezo input on the PDH box, however, the lock breaks and I cannot reacquire it.  We are using an SR650 to add in the signal from the network analyzer and that has worked. We also swapped the 20 dB attenuator for a box which mimics the boost functionality (-20 dB above 100 Hz, 0 dB below 6Hz).  I took some spectra with the SR750, and will get some more with the network analyzer once Alberto has finished with it. 

The SR750 spectra is posted below.  The SR750 only goes up to 100 kHz, so I will have to use the network analyzer to get the full range. 

  2702   Tue Mar 23 15:38:26 2010 AlbertoUpdateelogelog just restarted

I found the elog down and I restarted it.

Then, after few seconds it was down again. Maybe someone else was messing with it. I restarted an other 5 times and eventually it came back up.

  2701   Tue Mar 23 10:11:13 2010 steveUpdatePEM small air cond unit failed yesterday

Old control room air condition failed yesterday around noon. It was blowing 80-85F hot air for about 2-3 hours at racks 1Y4-7 and the entry room 103

  2700   Tue Mar 23 09:55:20 2010 KojiUpdateIOOvac envelope has to be sealed as antproof for overnight

Roger.

Quote:

 This seal is good for daily use- operation only. The IFO has to be sealed  with light metal doors every night so ants and other bugs can not find their way in.

 

  2699   Tue Mar 23 09:37:36 2010 steveUpdateIOOvac envelope has to be sealed as antproof for overnight

Quote:

This is the first touch to the MC mirrors after the earthquake on 16th.

  • I made an aluminum access connector so that we can work on the MC even the door is open. We still can be able to open the aluminum tube. The photos are attached. Steve, could you please look it at a glance whether the seal is enough or not.
  • MC resonances were flashing. Align MC2 and MC3 so that we have many TEM00s.
  • Found c1vmesus2 gone mad. Restarted remotely according to the wiki entry. 
  • Reset the MC coil output matrix to 1. (Previously, balance was adjusted so that A2L was minimized.)
  • Excite MC2 Pitch/Yaw at 8 and 9 Hz, looking at the peaks in the MC-MCL output. Move MC2 Pitch/Yaw so that the peak
    is reduced. (*)
  • MC1/MC3 were aligned so that we get the maximum transmission (or minimum reflection). (**)
  • Repeat (*) and (**)

So far, I have aligned in Yaw such that the yaw peak is minimized.

 This seal is good for daily use- operation only. The IFO has to be sealed  with light metal doors every night so ants and other bugs can not find their way in.

Our janitor Kevin is mopping the hole IFO room floor area with 5%  ant killing solution in water in order to discourage bugs getting close to our openings of the vented chamber.

You may be sensitive to this chemical too.  Do not open chamber till after lunch.

  2698   Tue Mar 23 00:31:51 2010 KojiUpdateIOOMC realigned

This is the first touch to the MC mirrors after the earthquake on 16th.

  • I made an aluminum access connector so that we can work on the MC even the door is open. We still can be able to open the aluminum tube. The photos are attached. Steve, could you please look it at a glance whether the seal is enough or not.
  • MC resonances were flashing. Align MC2 and MC3 so that we have many TEM00s.
  • Found c1vmesus2 gone mad. Restarted remotely according to the wiki entry. 
  • Reset the MC coil output matrix to 1. (Previously, balance was adjusted so that A2L was minimized.)
  • Excite MC2 Pitch/Yaw at 8 and 9 Hz, looking at the peaks in the MC-MCL output. Move MC2 Pitch/Yaw so that the peak
    is reduced. (*)
  • MC1/MC3 were aligned so that we get the maximum transmission (or minimum reflection). (**)
  • Repeat (*) and (**)

So far, I have aligned in Yaw such that the yaw peak is minimized.

  2697   Mon Mar 22 23:37:32 2010 MottUpdateABSLPLL reconstructed

 

We have managed to lock the PLL to reasonable stability. The RF input is attenuated by 26 dBm and the beat signal locks very close to the carrier of the marconi (the steps on the markers of the spectrum analyzer are coarse).  We can use the marconi and the local boost of the pdh box to catch the lock at 0 gain.  Once the lock is on, the gain can be increased to stabilize the lock.  The locked signals are shown in the first photo (the yellow is the output of the mixer and the blue is the output to the fast input of the laser.  If the gain is increased too high, the error signal enters an oscillatory regime, which probably indicates we are overloading the piezo.  This is shown in the second photo, the gain is being increased in time and we enter the non-constant regime around mid-way through.

Tomorrow I will use this locked system to measure the PZT response (finally!).

  2696   Mon Mar 22 22:11:26 2010 MottUpdateABSLPLL reconstructed

 

 It looks like the PLL drifted alot over the weekend, and we couldn't get it back to 9 dBm.  We switched back to the new focus wideband PD to make it easier to find the beat signal.  I replaced all the electronics with the newly fixed UPDH box (#17) and we aligned it to the biggest beat frequency we could get, which ended up being -27 dBm with a -6.3V DC signal from the PD.  

Locking was still elusive, so we calculated the loop gain and found the UGF is about 45 kHz, which is too high.  We added a 20 dB attenuator to the RF input to suppress the gain and we think we may have locked at 0 gain.  I am going to add another attenuator (~6 dB) so that we can tune the gain using the gain knob on the UPDH box.  

Finally, attached is a picture of the cable that served as the smb - BNC adaptor for the DC output of the PD.  The signal was dependent on the angle of the cable into the scope or multimeter.  It has been destroyed so that it can never harm another innocent experiment again!

  2695   Mon Mar 22 16:57:45 2010 josephb,daisuke, alexConfigurationComputersMegatron test points working again

We changed the pointer on /cvs/cds/caltech/target/gds/bin/awgtpman from

/opt/gds/awgtpman    to

/cvs/cds/caltech/target/gds/bin/awgtpman.091215.

Then killed the megatron framebuilder and testpoint manager (daqd, awgtpman), restarted, hit the daq reload button from the GDS_TP screen. 

This did not fix everything. However, it did seem to fix the problem where it needed a rtl_epics under the root directory which did not exist.  Alex continued to poke around.  When next he spoke, he claimed to have found a problem in the daqdrc file.  Specifically, the cvs/cds/caltech/target/fb/ daqdrc file.

set gds_server = "megatron" "megatron" 10 "megatron" 11;

He said this need to be:

set gds_server = "megatron" "megatron"  11 "megatron" 12;


However, during this, I had looked file, and found dataviewer working, while still with the 10 and 11.  Doing a diff on a backup of daqdrc, shows that Alex also changed

set controller_dcu=10  to set controller_dcu=12, and commented the previous line. 

He also changed set debug=2 to set debug=0.

In a quick test, we changed the 11 and 12 back to 10 and 11, and everything seemed to work fine.  So I'm not sure what that line actually does.  However, the set controller_dcu line seems to be important, and probably needs  to be set to the dcu id of an actually running module (it probably doesn't matter which one, but at least one that is up).  Anyways, I set the gds_server line back to 11 and 12, just in case there's numerology going on.

I'll add this information to the wiki.

  2694   Mon Mar 22 11:37:09 2010 steveUpdatePEMjackhammering

There was more jackhammering this morning just about 20 ft north-west of the beamsplitter chamber, outside.

  2693   Mon Mar 22 10:07:30 2010 KojiSummaryElectronicsUPDH Box #17: Ready

For your reference: Voltage noise of LM7815/LM7915 (with no load)

  2692   Mon Mar 22 02:03:57 2010 ranaSummaryElectronicsUPDH Box #17: Ready

It took too long to get this box ready for action. I implemented all of the changes that I made on the previous one (#1437). In addition, since this one is to be used for phase locking, I also made it have a ~flat transfer function. With the Boost ON, the TF magnitude will go up like 1/f below ~1 kHz.

The main trouble that I had was with the -12V regulator. The output noise level was ~500 nV/rHz, but there was a large oscillation at its output at ~65 kHz. This was showing up in the output noise spectrum of U1 (the first op-amp after the mixer). Since the PSRR of the OP27 is only ~40 dB at such a high frequency, it is not strange to see the power supply noise showing up (the input referred noise of the OP27 is 3.5 nV/rHz, so any PS noise above ~350 nV/rHz becomes relavent).

I was able to tame this by putting a 10 uF tantalum cap on the output of the regulator. However, when I replaced the regulator with a LM7912 from the blue box, it showed an output noise that went up like 1/f below 50 kHz !! I replaced it a couple more times with no benefit. It seems that something on the board must now be damaged. I checked another of the UPDH boxes, and it has the same high frequency oscillation but not so much excess voltage noise. I found that removing the protection diode on the output of the regulator decreased the noise by a factor of ~2. I also tried replacing all of the 1 uF caps that are around the regulator. No luck.

Both of the +12 V regulators seem fine: normal noise levels of ~200 nV/rHz and no oscillations.

Its clear that the regulator is not functioning well and my only guess is that its a layout issue on the board or else there's a busted component somewhere that I can't find. In any case, it seems to be functioning now and can be used for the phase locking and PZT response measurements.

  2691   Sun Mar 21 21:02:39 2010 KojiUpdatePSLEOM waist size
You don't need a lengthy code for this. It is obvious that the spot size at the distance L is minimum when L =
zR, where zR is the Rayleigh range. That's all.

Then compare the spot size and the aperture size whether it is enough or not.

It is not your case, but if the damage is the matter, just escape to the large zR side. If that is not possible
because of the aperture size, your EOM is not adequate for your purpose.
  2690   Sun Mar 21 20:08:20 2010 kiwamu, ranaUpdatePSLEOM wasit size

We are going to set the waist size to 0.1 mm for the beam going through the triple resonant EOM on a new PSL setup.

When we were drawing a new PSL diagram, we just needed to know the waist size at the EOM in order to think about mode matching.

waist.png

This figure shows the relation between the waist size and the spot size at the aperture of the EOM.

The x-axis is the waist size, the y-axis is the spot size. It is clear that there is a big clearance at 0.1 mm waist size. This is good.

Also it is good because the waist size is much above the damage threshold of the EO crystal (assuming 1W input).

The attached file is the python code for making this plot.

  2689   Sun Mar 21 19:25:29 2010 ranaUpdatePEMGuralp2 centered again

From this morning, now in calibrated units, and with the Güralp self noise spec from the Güralp manual.

  2688   Sat Mar 20 18:34:19 2010 kiwamuSummaryElectronicsRE:advantege of our triple resonant EOM

Yes, I found it.

Their advantage is that their circuit is isolated at DC because of the input capacitor.

And it is interesting that the performance of the circuit in terms of gain is supposed to be roughly the same as our transformer configuration.

  2687   Fri Mar 19 23:03:41 2010 ranaUpdatePEMGuralp2 centered again

I went and double-checked and aligned the styrofoam cooler at ~5:00 UTC. It was fine, but we really need a better huddling box. Where's that granite anyway?

Here's the new Huddle Test output. This time I show the X-axis since there's some coherence now below 0.1 Hz.

You'll also notice that the Wiener filter is now beating the FD subtraction. This happened when I increased the # of taps to 8000. Looks like the noise keeps getting lower as I increase the number of taps, but this is really a kind of cheat if you think about it carefully.

  2686   Fri Mar 19 21:15:33 2010 ranaUpdatePEMGuralp2 centered again

This trend of the last 200 days shows that GUR2 has been bad forever...until now anyways.

  2685   Fri Mar 19 18:00:14 2010 jenneUpdatePEMGuralp2 centered again

[Jenne, Sanjit]

It looks like Steve used a GND-12V supply to power the Guralp through the little breakout box (the box is for checking the centering of the mass).  This is BAD.  The Guralps want +/- 12V.

We centered all of the channels on Gur2, and checked the channels on Gur1, so we'll see how they're feeling after a while.

  2684   Thu Mar 18 21:42:26 2010 KojiUpdateABSLPLL reconstructed

I checked the setup further more.

  • I replaced the PD from NewFocus 1GHz one to Thorlabs PDA255.
  • I macthed the power level of the each beam.

Now I have significant fraction of beating (30%) and have huge amplitude (~9dBm).
The PLL can be much more stable now.

Koji

  2683   Thu Mar 18 19:00:04 2010 KojiSummaryElectronicsadvantege of our triple resonant EOM

Did you find what is the merit of their impedance matching technique?

Quote:

In this LVC meeting I discussed about triple resonant EOMs with Volker who was a main person of development of triple resonant EOMs at University of Florida.

Actually his EOM had been already installed at the sites. But the technique to make a triple resonance is different from ours.

They applied three electrodes onto a crystal instead of one as our EOM, and put three different frequencies on each electrode.

For our EOM, we put three frequencies on one electrode. You can see the difference in the attached figure. The left figure represents our EOM and the right is Volker's.

Then the question is; which can achieve better modulation efficiency ?

Volker and I talked about it and maybe found an answer,

 We believe our EOM can be potentially better because we use full length of the EO crystal.

This is based on the fact that the modulation depth is proportional to the length where a voltage is applied onto.

The people in University of Florida just used one of three separated parts of the crystal for each frequency.

 

  2682   Thu Mar 18 15:33:17 2010 kiwamuSummaryElectronicsadvantege of our triple resonant EOM

In this LVC meeting I discussed about triple resonant EOMs with Volker who was a main person of development of triple resonant EOMs at University of Florida.

Actually his EOM had been already installed at the sites. But the technique to make a triple resonance is different from ours.

They applied three electrodes onto a crystal instead of one as our EOM, and put three different frequencies on each electrode.

For our EOM, we put three frequencies on one electrode. You can see the difference in the attached figure. The left figure represents our EOM and the right is Volker's.

Then the question is; which can achieve better modulation efficiency ?

Volker and I talked about it and maybe found an answer,

 We believe our EOM can be potentially better because we use full length of the EO crystal.

This is based on the fact that the modulation depth is proportional to the length where a voltage is applied onto.

The people in University of Florida just used one of three separated parts of the crystal for each frequency.

  2681   Thu Mar 18 13:40:35 2010 KojiUpdateABSLPLL reconstructed

We use the current PLL just now, but the renewal of the components are not immediate as it will take some time. Even so we need steady steps towards the better PLL. I appreciate your taking care of it.

Quote:

Quote:

Last night (Mar 17) I checked the PLL setup as Mott had some difficulty to get a clean lock of the PLL setting.

  • I firstly found that the NPRO beam is not going through the Faraday isolator well. This was fixed by aligning the steering mirrors before the Faraday.
     
  • The signal from the RF PD was send to the RF spectrum analyzer through a power splitter. This is a waist of the signal. It was replaced to a directional coupler.
  • Tee-ing the PZT feedback to the oscilloscope was producing the noise in the laser frequency. I put the oscilloscope to the 600Ohm output of the SR560, while connectiong the PZT output to the 50Ohm output.
  • In addition, 6dB+6dB attenuators have been added to the PZT feedback signal.

Now the beating signal is much cleaner and behave straight forward. I will add some numbers such as the PD DC output, RF levels, SR560 settings...

Now I am feeling that we definitely need the development of really clean PLL system as we use PLL everywhere! (i.e. wideband PD, nice electronics, summing amplifiers, stop poking SR560, customize/specialize PDH box, ...etc)

I also had noticed the progressive change of the aux NPRO alignment to the Farady.

I strongly agree about the need of a good and robust PLL.

By modifying the old PDH box (version 2008) eventually I was able to get a PLL robust enough for my purposes. At some point that wasn't good enough for me either.

I then decided to redisign it from scratch. I'm going to work on it. Also because of my other commitments, I'd need a few days/1 week for that. But I'd still like to take care of it. Is it more urgent than that?

 

  2680   Thu Mar 18 12:27:56 2010 AlbertoUpdateABSLPLL reconstructed

Quote:

Last night (Mar 17) I checked the PLL setup as Mott had some difficulty to get a clean lock of the PLL setting.

  • I firstly found that the NPRO beam is not going through the Faraday isolator well. This was fixed by aligning the steering mirrors before the Faraday.
     
  • The signal from the RF PD was send to the RF spectrum analyzer through a power splitter. This is a waist of the signal. It was replaced to a directional coupler.
  • Tee-ing the PZT feedback to the oscilloscope was producing the noise in the laser frequency. I put the oscilloscope to the 600Ohm output of the SR560, while connectiong the PZT output to the 50Ohm output.
  • In addition, 6dB+6dB attenuators have been added to the PZT feedback signal.

Now the beating signal is much cleaner and behave straight forward. I will add some numbers such as the PD DC output, RF levels, SR560 settings...

Now I am feeling that we definitely need the development of really clean PLL system as we use PLL everywhere! (i.e. wideband PD, nice electronics, summing amplifiers, stop poking SR560, customize/specialize PDH box, ...etc)

I also had noticed the progressive change of the aux NPRO alignment to the Farady.

I strongly agree about the need of a good and robust PLL.

By modifying the old PDH box (version 2008) eventually I was able to get a PLL robust enough for my purposes. At some point that wasn't good enough for me either.

I then decided to redisign it from scratch. I'm going to work on it. Also because of my other commitments, I'd need a few days/1 week for that. But I'd still like to take care of it. Is it more urgent than that?

  2679   Thu Mar 18 10:46:51 2010 KojiUpdateABSLPLL reconstructed

Last night (Mar 17) I checked the PLL setup as Mott have had some difficulty to get a clean lock of the PLL setting.

  • I firstly found that the NPRO beam is not going through the Faraday isolator well. This was fixed by aligning the steering mirrors before the Faraday.
     
  • The signal from the RF PD was send to the RF spectrum analyzer through a power splitter. This is a waist of the signal. It was replaced to a directional coupler.
     
  • Tee-ing the PZT feedback to the oscilloscope was producing the noise in the laser frequency. I put the oscilloscope to the 600Ohm output of the SR560, while connectiong the PZT output to the 50Ohm output.
     
  • In addition, 6dB+6dB attenuators have been added to the PZT feedback signal.

Now the beating signal is much cleaner and behave straight forward. I will add some numbers such as the PD DC output, RF levels, SR560 settings...

Now I am feeling that we definitely need the development of really clean PLL system as we use PLL everywhere! (i.e. wideband PD, nice electronics, summing amplifiers, stop poking SR560, customize/specialize PDH box, ...etc)

  2678   Thu Mar 18 08:49:51 2010 steveUpdatePEMGuralp2 centered again ?

 

Someone adjusted the Guralp2 mass position last night??

NO

  2677   Tue Mar 16 09:37:30 2010 steveUpdateSUSeq 4.4 seen by oplevs and osems

The oplev plots clearly show the alignment effect of this eq.

  2676   Tue Mar 16 05:10:39 2010 kojiUpdateEnvironmentEarthquake, Mag 4.4

Some of the suspensions got watchdog tripped -> enabled -> damped.

The MC mirrors got slightly misaligned.

 

  2675   Tue Mar 16 04:22:02 2010 JenneUpdateEnvironmentEarthquake, Mag 4.4

Earthquake Details

Magnitude 4.4
Date-Time
  • Tuesday, March 16, 2010 at 11:04:00 UTC
  • Tuesday, March 16, 2010 at 04:04:00 AM at epicenter
Location 33.998°N, 118.072°W
Depth 18 km (11.2 miles)
Region GREATER LOS ANGELES AREA, CALIFORNIA
Distances
  • 2 km (1 miles) ENE (58°) from Pico Rivera, CA
  • 4 km (3 miles) ESE (115°) from Montebello, CA
  • 4 km (3 miles) SSW (207°) from Whittier Narrows Rec. Area, CA
  • 9 km (5 miles) NE (39°) from Downey, CA
  • 9 km (6 miles) SSW (206°) from El Monte, CA
  • 17 km (11 miles) ESE (112°) from Los Angeles Civic Center, CA
Location Uncertainty horizontal +/- 0.4 km (0.2 miles); depth +/- 0.9 km (0.6 miles)
Parameters Nph=122, Dmin=5 km, Rmss=0.37 sec, Gp= 25°,
M-type=local magnitude (ML), Version=2
Source
Event ID

ci14601172

-----------------------------------------

MC1 and MC3 seem to have kept themselves together, but all the other optics' watchdogs tripped.

  2674   Mon Mar 15 16:39:36 2010 steveUpdatePEMGuralp2 centered

Quote:

Untitled.png

 

Guralp 2 centered.The mass position offsets are: E-W 0.05V, N-S 0V, Z 0.4V

Guralp 1: E-W -0.1V, N-S -0.25V, Z 0V measured, not adjusted

The GUR2_X channel has an offset. See plot below when seismometers are disconnected. This offset has to be removed.

NOTE: this huddle is on bad-soft ground-lenoleum tile from prehistoric Flintstone age

 

  2673   Mon Mar 15 09:43:47 2010 steveUpdatePEMmore sandblasting today

Do not open IFO vacuum envelope today! They are sandblasting again at CES

  2672   Sat Mar 13 22:16:03 2010 ranaUpdateWienerFilteringGuralp Huddle Test software

I used some recent better data to try for better Z subtraction.

Dmass helped me understand that sqrt(1-Coherence) is a good estimate of the theoretical best noise subtraction residual. This should be added to DTT. For reference the Jan statistic is the inverse of this.

This should get better once Steve centers the Guralps. 

  2671   Sat Mar 13 21:30:22 2010 ranaUpdatePEMStyrofoam Cooler on the Seismos

Untitled.png

 

  2670   Fri Mar 12 17:08:22 2010 steveConfigurationVACbg-RGAscan at d18

RGA scan of rga-region only at day 18   This is the back ground of the rga with some calibration gas.

  2669   Fri Mar 12 13:52:18 2010 ZachUpdateelogelog restarted

 The elog was down and I ran the restart script.

  2668   Thu Mar 11 17:51:04 2010 KojiUpdateSUSRecent status of SOSs

Jenne, Koji

Recent status of SOSs:

We completed one of the suspension (ITMY).

ITMX: 6 Magnets, standoffs, and guide rod glued / balance to be confirmed / needs to be baked
ITMY: 6 Magnets, standoffs, and guide rod glued / balance confirmed / needs to be baked
SRM: 6 Magnets, one standoff, and guide rod glued,  / waiting for the release from the gluing fixture.
PRM: one standoff, and guide rod glued / waiting for the magnet gluing. 

We think we solved all the problems for hanging the suspensions.

--- Magnet gluing fixture ---

  • There is the two kinds of fixtures. Neither does work propery in the original form!
  • The height of the side magnets should be finely adjusted by changing the teflon sheets beneath the optics in the fixture.
  • Be aware of the polarity of the fixture in terms of the side magnets
  • Wrongly glued magnets (and others) can be removed by a razor blade with some amount of acetone.
  • The pickle picker frequently knocks the magnets down during the release. Don't s be down in the dumps too much.

--- Suspending the mirror ---

  • The wire winches must be carefully attached to the suspension tower such that the wires are not streached during fastening the clamps.
  • There are a couple variations of the drawings for SOS. The one we have has #4-40 for the earthquake stops at the bottom.
    Zach and Mott made the EQ stops with the right size.
  2667   Thu Mar 11 15:24:51 2010 steveMetaphysicsEnvironmentIFO was well protected

Quote:

There is a planned power outage tomorrow, Saturday from 7am till midnight.

I vented all annulies and switched to ALL OFF configuration. The small region of the RGA is still under vacuum.

The vac-rack: gauges, c1vac1 and UPS turned off.

 It turns out that we perfecly timed the big one

In the process of finding the signal of the big chilean earthquake I just realized that we were all off

  2666   Wed Mar 10 15:02:58 2010 ranaUpdatePEMStyrofoam Cooler on the Seismos

This is the spectra and coherence from a quiet time last night. I've lowered the Guralp cal by a factor of 2 to account for the fact that the gain in the breakout box is actually 20 and not 10 as I previously said.

The AD620 stage in the front part has a gain of 10 and then there's a single-to-differential stage in the output which gives us a gain of 2. The DTT cal in counts is now 3.8e-9 (m/s)/count.

The second plot shows the Guralp and Ranger signals at the ADC input (converted from counts to Volts for usefulness). The thick grey line is the expected noise of the Guralp breakout box

(mainly the AD620) propagated to the ADC (via multiplication by 2). It looks like the preamp board should not be a problem as long as we can reach the AD620 limit.

So the excess noise in the Guralp is not the fault of the preamp, but more likely the mounting and insulation of the seismometers.

  2665   Tue Mar 9 12:06:53 2010 rana, JenneUpdatePEMStyrofoam Cooler on the Seismos

Looks like the GUR2_X signal is bad. Jenne says that we need to center it mechanically before the signals will become useful again. Maybe Steve will do this - instructions are in the manuel ?

  2664   Tue Mar 9 09:32:31 2010 KojiSummaryGeneralWideband measurement of Fast PZT response

I have measured a wideband response of the fast PZT in the LWE NPRO 700mW in the Alberto's setup.
This is a basic measurement to determine how much phase modulation we can obtain by actuating the fast PZT,
primarily for the green locking experiment.

RESULT

  • Above 200kHz, there are many resonances that screws up the phase.
     
  • Modulation of 0.1rad can be easily obtained even at 10MHz if the modulation frequency is scanned.
     
  • Change of the laser frequency in DC was observed depending on the modulation frequency.
    i.e. At the resonance the laser frequency escaped from the RF spectrum analyzer.
    This may induced by the heat dissipation in the PZT causing the temperature change of the crystal.
     
  • Some concerns: Is there any undesired AM by the PZT modulation?

---

METHOD

1. Locked the PLL of for the PSL-NPRO beating at 20MHz.

2. Added the modulation signal to the NPRO PZT input.
I used the output of the network analyzer sweeping from 100kHz to 1MHz.

3. Measured the transfer function from the modulation input to the PLL error signal.
The PLL error is sensitive to the phase fluctuation of the laser. Found that the first resonance is at 200kHz.
The TF is not valid below 3kHz where the PLL suppresses the modulation.

4. Single frequency modulation: Disconnected the PLL setup.
Plug Marconi into the fast PZT input and modulate it at various frequencies.
Observing with the RF spectrum analyzer, I could see strong modulation below 1MHz.
It turned out later that the TF measurement missed the narrow peaks of the resonances due to the poor freq resolution.

Also the modulation depth varies frequency by frequency because of the resonances.
Scanned the frequency to have local maximum of the modulation depth. Adjusted the
modulation amplitude such that the carrier is suppressed
(J0(m)=0 i.e. m~2.4). As I could not obtain
the carrier suppression at above 1MHz, the height of the carrier and the sidebands were measured.

The modulation frequency was swept from 100kHz to 10MHz.

5. Calibration. The TF measured has been calibrated using the modulation depth obtained at 100Hz,
where the resonance does not affect the response yet.

The responce of the PZT was ~10MHz/V below 30kHz. Looks not so strange although this valure is
little bit high from the spec (2MHz/V), and still higher than my previous experience at TAMA (5MHz/V).
Note that this calibration does not effect to the modulation depth of the single freq measurement as they are independent.

ELOG V3.1.3-